# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 2.2 TW Curvature Tensors
[~$[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\mathcal{V}^{\alpha}\equiv\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{V}^{\beta}~{}~{}~{},~{}~{}~{}[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\mathcal{V}_{\beta}\equiv-\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{V}_{\alpha}$ (2.18)
@@
$\mathcal{K}^{\alpha}{}_{\beta\mu\nu}=$ $\partial_{\mu}\tilde{\Gamma}^{\alpha}{}_{\beta\nu}-\partial_{\nu}\tilde{\Gamma}^{\alpha}{}_{\beta\mu}+\tilde{\Gamma}^{\rho}_{\ \beta\nu}\tilde{\Gamma}^{\alpha}{}_{\mu\rho}-\tilde{\Gamma}^{\rho}{}_{\beta\mu}\tilde{\Gamma}^{\alpha}{}_{\nu\rho}~{}~{}~{}.$ (2.19)
@@
$\mathcal{R}^{a}_{\ bcd}=\partial_{c}\Pi^{a}_{\ bd}-\partial_{d}\Pi^{a}_{\ bc}+\Pi^{e}_{\ bd}\Pi^{a}_{\ ce}-\Pi^{e}_{\ bc}\Pi^{a}_{\ de}$ (2.20)
@@
$R^{a}_{\ bcd}=\partial_{c}\Gamma^{a}_{\ bd}-\partial_{d}\Gamma^{a}_{\ bc}+\Gamma^{e}_{\ bd}\Gamma^{a}_{\ ce}-\Gamma^{e}_{\ bc}\Gamma^{a}_{\ de}~{}~{}~{}.$ (2.21)
@@
${\mathcal{K}}^{a}{}_{bcd}=-{\mathcal{K}}^{a}{}_{bdc}~{}~{}~{},~{}~{}~{}R^{a}{}_{bcd}=-R^{a}{}_{bdc}~{}~{}~{},~{}~{}~{}\mathcal{R}^{a}{}_{bcd}=-{\mathcal{R}}^{a}{}_{bdc}~{}~{}~{}.$ (2.22)
@@
$\mathcal{R}_{bd}\equiv$ $\mathcal{R}^{a}{}_{bad}$ (2.23)
$R_{bd}\equiv$ $R^{a}{}_{bad}$ (2.24)
$\mathcal{K}_{\beta\nu}=$ $\mathcal{K}^{\alpha}{}_{\beta\alpha\nu}~{}~{}~{}.$ (2.25)
It should be noted that both $\mathcal{R}_{bd}$ and $\mathcal{K}_{\beta\nu}$ are symmetric while $R_{bd}$ is not. The antisymmetric part of $R_{bd}$ is proportional to the curl of $\alpha_{b}$ as well as the antisymmetric part of $\mathcal{P}_{bd}$
$\mathcal{P}_{[bd]}=\frac{1}{d+1}R_{[bd]}=-\partial_{[b}\alpha_{d]}~{}~{}~{}.$ (2.26)
It is notable that the only non-vanishing components of $\mathcal{K}_{\beta\nu}$ are those on the manifold $\mathcal{M}$, namely $\mathcal{K}_{bn}$. To form scalars such as the Ricci tensor requires a metric, to which we turn to next.
2.3 The TW Action
@@
$G_{\alpha\beta}$ $=\delta^{a}_{\,\,\alpha}\delta^{b}_{\,\,\beta}\,g_{ab}-\lambda_{0}^{2}g_{\alpha}g_{\beta}$ (2.27)
$G^{\alpha\beta}$ $=g^{ab}(\delta^{\alpha}_{\ a}-g_{a}\Upsilon^{\alpha})(\delta^{\beta}_{\ b}-g_{b}\Upsilon^{\alpha})-\lambda_{0}^{-2}\Upsilon^{\alpha}\Upsilon^{\beta}~{}~{}~{}.$ (2.28)
@@
$g_{a}=$ $-\frac{1}{d+1}\partial_{a}\log\sqrt{|g|}=-\tfrac{1}{{\rm d}+1}\hat{\Gamma}^{c}{}_{ac}$ (2.30)
$\hat{\Gamma}^{c}{}_{ab}=$ $\frac{1}{2}g^{cm}\left(\partial_{(a}g_{b)m}-\partial_{m}g_{ab}\right)~{}~{}~{}.$ (2.31)
@@
Using Eq. (2.31) we can expand the Riemann curvature tensor as
$R^{a}_{\ bcd}=$ $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}$ (2.33)
$Q^{a}_{\ bcd}=$$\hat{\nabla}_{c}C^{a}_{\ bd}-\hat{\nabla}_{d}C^{a}_{\ bc}+C^{e}_{\ bd}C^{a}_{\ ce}-C^{e}_{\ bc}C^{a}_{\ de}$ (2.34)
where $\hat{R}^{a}_{\ bcd}$ is the usual Levi-Civita Riemann tensor, built from $\hat{\Gamma}^{a}{}_{bc}$ rather than $\Gamma^{a}{}_{bc}$ as in Eq. (2.21), and ^$\hat{\nabla}_{a}$ the Levi-Cita covariant derivative for the connection $\hat{\Gamma}^{c}{}_{ab}$. The Levi-Civita Ricci curvature and scalar along with the analogous versions involving $Q^{a}{}_{bcd}$ are defined in the usual way
$\hat{R}_{bd}=$ $\hat{R}^{a}{}_{bad}~{}~{}~{},~{}~{}~{}\hat{R}=g^{bd}\hat{R}_{bd}~{}~{}~{}$ (2.35)
$Q_{bd}=$ $Q^{a}{}_{bad}~{}~{}~{},~{}~{}~{}Q=g^{bd}Q_{bd}~{}~{}~{}.$ (2.36)
With the metric $G_{\alpha\beta}$ now in hand from Eq. (2.27), we can construct invariants on the Thomas cone such as
$\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}_{\alpha}{}^{\beta\mu\nu}\equiv$ $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}^{\rho}{}_{\sigma\kappa\tau}G_{\alpha\rho}G^{\beta\sigma}G^{\mu\kappa}G^{\nu\tau}$ (2.37)
$\mathcal{K}_{\alpha\beta}\mathcal{K}^{\alpha\beta}\equiv$ $\mathcal{K}^{\rho}{}_{\alpha\rho\beta}\mathcal{K}^{\sigma}{}_{\mu\sigma\nu}G^{\alpha\mu}G^{\beta\nu}$ (2.38)
$\mathcal{K}\equiv$ $\mathcal{K}_{\alpha\beta}G^{\alpha\beta}~{}~{}~{}.$ (2.39)
@@
$K_{bcd}\equiv$ $g_{\alpha}\mathcal{K}^{\alpha}{}_{bcd}$ (2.40)
$=$ ˘$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}$ (2.41)
$=$ $\nabla_{[c}\mathcal{P}_{d]b}+\text{\textcent}_{a}\mathcal{K}^{a}{}_{bcd}~{}~{}~{}$ (2.42)
where ˘$\breve{\nabla}_{c}{\mathcal{D}}_{db}$ is defined as in Eq. (2.16) and ¢$\text{\textcent}_{a}$ is defined as
¢$\text{\textcent}_{a}\equiv$ $g_{a}-\alpha_{a}=\tfrac{1}{{\rm d}+1}C^{c}{}_{ac}~{}~{}~{}.$ (2.43)
@@
$K_{bcd}=-K_{bdc}~{}~{}~{}.$ (2.44)
The projective Cotton-York tensor $K_{bcd}$ is clearly a tensor on both $\mathcal{M}$ and $\mathcal{N}$ as well as projectively invariant. In Eq. (2.40), the first line demonstrates the tensor character on $\mathcal{N}$ as recall both $g_{\alpha}$ and $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}$ are tensors on $\mathcal{N}$. The second line demonstrates projective invariance even when restricted to the manifold $\mathcal{M}$ as each of $\mathcal{D}_{ab}$, $g_{a}$, and $\mathcal{K}^{a}{}_{bcd}$ are projectively invariant. The third line demonstrates the tensor character on the manifold $\mathcal{M}$ as each of $\mathcal{P}_{ab}$, ¢$\text{\textcent}_{a}$, and $\mathcal{K}^{a}{}_{bcd}$ are tensors on $\mathcal{M}$. None of the lines alone demonstrate both tensorality and projective invariance on the manifold $\mathcal{M}$, once again showing the importance of defining the tensorial and projectively invariant ${\bar{{\mathcal{P}}}}_{ab}$ and ¯$\bar{\nabla}_{a}$ that we will introduce later.
The TW action consists of two components, the projective Einstein-Hilbert action $S_{\text{PEH}}$ and the projective Gauss-Bonnet action $S_{\text{PGB}}$, which are given by
@@
and
$S_{\text{PGB}}=-\frac{\tilde{J}_{0}c}{\lambda_{0}}\int d^{{\rm d}}xd\lambda\sqrt{|G|}\ \left(\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}_{\alpha}{}^{\beta\mu\nu}-4\mathcal{K}_{\alpha\beta}\mathcal{K}^{\alpha\beta}+\mathcal{K}^{2}\right)$ (2.46)
@@
Our goal in this paper is to maintain this manifest covariance and projective invariance after reducing to the ${\rm d}$-dimensional manifold $\mathcal{M}$. To illuminate the behavior on the spacetime manifold $\mathcal{M}$ we can expand $S$ and integrate out $\lambda$
$S=$ $-\frac{1}{2\kappa_{0}}\int d^{\rm d}x\sqrt{|g|}\mathcal{K}+J_{0}c\lambda_{0}^{2}\int d^{\rm d}x\sqrt{|g|}K_{bcd}K^{bcd}$ (2.47)
$-J_{0}c\int d^{\rm d}x\sqrt{|g|}\left(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2}\right)~{}~{}~{}.$ (2.48)
The rescaled coupling constants are
@@
On the manifold $\mathcal{M}$ we use the notation that indices are raised and lowered with $g^{ab}$ and $g_{ab}$, respectively
${\mathcal{K}}_{abcd}\equiv$ $g_{ma}{\mathcal{K}}^{m}{}_{bcd}$ (2.50)
${\mathcal{K}}^{abcd}=$ $g^{bp}g^{cm}g^{dn}{\mathcal{K}}^{a}{}_{pmn}$ (2.51)
$K^{bcd}\equiv$ $g^{bp}g^{cm}g^{dn}K_{pmn}~{}~{}~{}.$ (2.52)
@@
On the manifold $\mathcal{N}$ whether indices are raised with $g^{ab}$ or $G^{\alpha\beta}$ is unambiguous for the projective Cotton-York tensor and the last three indices of the projective Riemann tensor $\mathcal{K}^{a}{}_{bcd}$. This is because $G^{ab}=g^{ab}$ and the lambda components of these tensor vanish aside from $\mathcal{K}^{\lambda}{}_{bcd}$. Care must be taken, however, to be clear what is meant by lowering the first index of $\mathcal{K}^{a}{}_{bcd}$. On the manifold $\mathcal{M}$, we no longer have the full $G_{\alpha\beta}$ as its $\lambda$ components and $\lambda$-dependence have been integrated out, and so raising and lowering is done via $g^{ab}$ and $g_{ab}$, respectively.
@@
Quadratic dynamics for the fields can be seen in the projective Cotton-York tensor term when expanded using Eq. (2.40). The curvature tensors contain interactions as can be seen upon expanding in terms of the ${\rm d}$-dimensional $\mathcal{D}_{ab}$ and $\Pi^{c}{}_{ab}$ or ${\mathcal{P}}_{ab}$ and $\Gamma^{c}{}_{ab}$ as
$\mathcal{K}^{a}{}_{bcd}=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{D}_{d]b}=R^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (2.53)
$\mathcal{K}_{mn}=$ $\mathcal{R}_{mn}+(d-1)\mathcal{D}_{mn}=R_{mn}+(d-1)\mathcal{P}_{mn}-d\mathcal{P}_{[mn]}$ (2.54)
$\mathcal{K}=$ $\mathcal{R}+(d-1)\mathcal{D}=R+(d-1)\mathcal{P}~{}~{}~{}$ (2.55)
$\mathcal{R}\equiv$ $g^{ab}\mathcal{R}_{ab}~{}~{}~{},~{}~{}~{}\mathcal{D}\equiv g^{ab}\mathcal{D}_{ab}~{}~{}~{},~{}~{}~{}R\equiv g^{ab}R_{ab}~{}~{}~{},~{}~{}~{}{\mathcal{P}}\equiv g^{ab}{\mathcal{P}}_{ab}~{}~{}~{}.$ (2.56)
@@
All other curvature components vanish aside from $\mathcal{K}^{\lambda}{}_{bcd}$ which can be solved for using the expressions of the projective Cotton-York tensor in Eq. (2.40). We see explicitly that TW gravity is not simultaneously manifestly both general coordinate invariant on $\mathcal{M}$ and projectively invariant when written in terms of these expansions. If the curvature tensors are expanded through the first equality in each of the above, they are manifestly projectively invariant, but not manifestly tensorial. If the second equality is used in each, they are no longer manifestly projectively invariant, but they are manifestly tensorial. The same is true for the projective Cotton-York tensor as was shown in Eq. (2.40). In the next section, we introduce the manifestly both tensorial and projectively invariant ${\bar{{\mathcal{P}}}}_{ab}$ and ¯$\bar{\nabla}_{a}$ and expand the above tensors in terms of these.
@@
We conclude this section with a brief aside on the the values of the physical constants of TW gravity that are consistent with physical phenomena, specifically early Universe inflation. For ${\rm d}=4$ the units are
$\begin{split}[J_{0}]=\frac{ML^{2}}{T}\;,\quad[\mathcal{D}_{ab}]&=[R_{ab}]=L^{-2}\;,\quad\left[\kappa_{0}\right]=\frac{T^{2}}{ML}~{}~{}~{}.\cr\end{split}$ (2.57)
@@
$M_{p}=\sqrt{\frac{\hbar c}{8\pi G}}$ (2.58)
@@
a recent analysis in [23] suggests that if $\lambda_{0}\lesssim 6\cdot 10^{6}M_{p}^{-1}$ the inflaton $\phi\propto{\mathcal{P}}$ appearing naturally in TW gravity would have provided early Universe inflation matching current inflationary cosmology data within the $95\%$ confidence level. Specifically, it was found that to produce $N=60$ efolds, the parameters would have to satisfy the constraints $\lambda_{0}^{2}\lesssim 10^{3}J_{0}\kappa_{0}$ and $10^{8}\lesssim J_{0}\lesssim 4\cdot 10^{10}$ all within the $95\%$ confidence level. Furthermore, a value of $\kappa_{0}\sim M_{p}^{-2}$ would result in Einstein-Hilbert gravity + cosmological constant at the end of inflation within the $95\%$ confidence level  [^4] The value of $\kappa_{0}$ was not discussed in this way in [23]. However, it is easy to see that the TW action reduces to $(-2\kappa_{0})^{-1}(1+8\kappa_{0}J_{0}\lambda_{0}^{-1}\phi)R$ plus cosmological constant at the end of inflation if $\phi\propto{\mathcal{P}}$ settles to a constant value, with the traceless parts of $\mathcal{P}_{ab}$ equal to zero and assuming a Levi-Civita connection. The analysis in [23] leads to $18\kappa_{0}J_{0}\lambda_{0}^{-1}\phi<<$ 1 at the end of $N=60$ efolds of inflation within the $95\%$ confidence level. Thus the TW action reduces to approximately $(-2\kappa_{0})^{-1}R$ plus cosmological constant at the end of inflation, where we identify $\kappa_{0}\sim M_{p}^{-2}$ to match general relativity.. Putting this all together results in the above limit on $\lambda_{0}$. This suggest that the characteristic projective scale defined by $\lambda_{0}$ is less than a few million Planck lengths if it is TW gravity itself that is responsible for inflation. The importance of these results is that it appears inflation is automatically is built into the geometry of TW gravity with no additional field needed. The above analysis for the limit on $\lambda_{0}$ relies on the assumption that TW gravity must precisely match Einstein-Hilbert gravity + cosmological constant at the end of inflation to then transition to radiation dominance. Whether or not that is true depends on the specific interactions of TW gravity with matter because matter acts like radiation in the early Universe. These interactions are an ongoing avenue of research [18, 19, 45, 20].
2.4 Manifestly Projective Invariant form of the TW Field Equations

### v3

@@ II.2 TW Curvature Tensors
[~$[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\mathcal{V}^{\alpha}\equiv$ $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{V}^{\beta}$ (21)
[~$[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\mathcal{V}_{\beta}\equiv$ $-\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{V}_{\alpha}$ (22)
@@
$\mathcal{K}^{\alpha}_{\ \beta\mu\nu}=$ $\partial_{\mu}\tilde{\Gamma}^{\alpha}_{\ \nu\beta}-\partial_{\nu}\tilde{\Gamma}^{\alpha}_{\ \beta\mu}+\tilde{\Gamma}^{\rho}_{\ \beta\nu}\tilde{\Gamma}^{\alpha}_{\ \mu\rho}-\tilde{\Gamma}^{\rho}_{\ \beta\mu}\tilde{\Gamma}^{\alpha}_{\ \nu\rho}~{}.$ (23)
@@
$\mathcal{R}^{a}_{\ bcd}=\partial_{c}\Pi^{a}_{\ bd}-\partial_{d}\Pi^{a}_{\ bc}+\Pi^{e}_{\ bd}\Pi^{a}_{\ ce}-\Pi^{e}_{\ bc}\Pi^{a}_{\ de}$ (24)
@@
$R^{a}_{\ bcd}=\partial_{c}\Gamma^{a}_{\ bd}-\partial_{d}\Gamma^{a}_{\ bc}+\Gamma^{e}_{\ bd}\Gamma^{a}_{\ ce}-\Gamma^{e}_{\ bc}\Gamma^{a}_{\ de}~{}.$ (25)
@@
${\mathcal{K}}^{a}_{\ bcd}=$ $-{\mathcal{K}}^{a}_{\ bdc}$ (26)
$R^{a}_{\ bcd}=$ $-R^{a}_{\ bdc}$ (27)
$\mathcal{R}^{a}_{\ bcd}=$ $-{\mathcal{R}}^{a}_{\ bdc}~{}.$ (28)
@@
$\mathcal{R}_{bd}\equiv$ $\mathcal{R}^{a}_{\ bad}$ (29)
$R_{bd}\equiv$ $R^{a}_{\ bad}$ (30)
$\mathcal{K}_{\beta\nu}=$ $\mathcal{K}^{\alpha}_{\ \beta\alpha\nu}={\mathcal{K}}^{a}_{\ \beta a\nu}~{}.$ (31)
Notice that it matters not whether the contraction forming ${\mathcal{K}}_{\beta\nu}$ includes the $\lambda$-direction or not as ${\mathcal{K}}^{\alpha}_{\ \beta\mu\nu}$ vanishes when any of its lower indices equals $\lambda$. Furthermore, the only non-vanishing components of $\mathcal{K}_{\beta\nu}$ are those on the manifold $\mathcal{M}$, namely $\mathcal{K}_{bn}$. It should be noted that both $\mathcal{R}_{bd}$ and $\mathcal{K}_{\beta\nu}$ are symmetric while $R_{bd}$ is not. The antisymmetric part of $R_{bd}$ is proportional to the curl of $\alpha_{b}$ as well as the antisymmetric part of $\mathcal{P}_{bd}$
$\mathcal{P}_{[bd]}=\frac{1}{d+1}R_{[bd]}=-\partial_{[b}\alpha_{d]}~{}$ (32)
where here and throughout the paper, square brackets denote antisymmetrization of enclosed indices as described in Eq. (260) of Appendix A. To form scalars such as the Ricci tensor requires a metric, to which we turn to next.
II.3 The TW Action
@@
$G_{\alpha\beta}$ $=\delta^{a}_{\,\,\alpha}\delta^{b}_{\,\,\beta}\,g_{ab}-\lambda_{0}^{2}g_{\alpha}g_{\beta}$ (33)
$G^{\alpha\beta}$ $=g^{ab}(\delta^{\alpha}_{\ a}-g_{a}\Upsilon^{\alpha})(\delta^{\beta}_{\ b}-g_{b}\Upsilon^{\alpha})-\lambda_{0}^{-2}\Upsilon^{\alpha}\Upsilon^{\beta}~{}.$ (34)
@@
$g_{a}=$ $-\tfrac{1}{d+1}\partial_{a}\log\sqrt{|g|}=-\tfrac{1}{{\rm d}+1}\hat{\Gamma}^{c}{}_{ac}$ (36)
$\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}\left(\partial_{(a}g_{b)m}-\partial_{m}g_{ab}\right)~{}.$ (37)
@@
Using Eq. (37) we can expand the Riemann curvature tensor as
$R^{a}_{\ bcd}=$ $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}$ (39)
$Q^{a}_{\ bcd}=$$\hat{\nabla}_{c}C^{a}_{\ bd}-\hat{\nabla}_{d}C^{a}_{\ bc}+C^{e}_{\ bd}C^{a}_{\ ce}-C^{e}_{\ bc}C^{a}_{\ de}$ (40)
where $\hat{R}^{a}_{\ bcd}$ is the usual Levi-Civita Riemann tensor, built from $\hat{\Gamma}^{a}{}_{bc}$ rather than $\Gamma^{a}{}_{bc}$ as in Eq. (25), and ^$\hat{\nabla}_{a}$ the Levi-Cita covariant derivative for the connection $\hat{\Gamma}^{c}{}_{ab}$. The Levi-Civita Ricci curvature and scalar along with the analogous versions involving $Q^{a}{}_{bcd}$ are defined in the usual way
$\hat{R}_{bd}=$ $\hat{R}^{a}{}_{bad}~{}~{}~{},~{}~{}~{}\hat{R}=g^{bd}\hat{R}_{bd}~{}~{}~{}$ (41)
$Q_{bd}=$ $Q^{a}{}_{bad}~{}~{}~{},~{}~{}~{}Q=g^{bd}Q_{bd}~{}.$ (42)
With the metric $G_{\alpha\beta}$ now in hand from Eq. (33), we can construct invariants on the Thomas cone such as
$\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}_{\alpha}{}^{\beta\mu\nu}\equiv$ $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}^{\rho}{}_{\sigma\kappa\tau}G_{\alpha\rho}G^{\beta\sigma}G^{\mu\kappa}G^{\nu\tau}$ (43)
$\mathcal{K}_{\alpha\beta}\mathcal{K}^{\alpha\beta}\equiv$ $\mathcal{K}^{\rho}{}_{\alpha\rho\beta}\mathcal{K}^{\sigma}{}_{\mu\sigma\nu}G^{\alpha\mu}G^{\beta\nu}$ (44)
$\mathcal{K}\equiv$ $\mathcal{K}_{\alpha\beta}G^{\alpha\beta}={\mathcal{K}}_{ab}g^{ab}~{}.$ (45)
@@
A portion of the expansion of Eq. (43) produces the projective Cotton-York tensor
$K_{bcd}\equiv$ $g_{\alpha}\mathcal{K}^{\alpha}{}_{bcd}$ (46)
$=$ ˘$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}$ (47)
$=$ $\nabla_{[c}\mathcal{P}_{d]b}+\text{\textcent}_{a}\mathcal{K}^{a}{}_{bcd}~{}~{}~{}$ (48)
where ˘$\breve{\nabla}_{c}{\mathcal{D}}_{db}$ is defined as in Eq. (19) and ¢$\text{\textcent}_{a}$ is defined as
¢$\text{\textcent}_{a}\equiv$ $g_{a}-\alpha_{a}=\tfrac{1}{{\rm d}+1}C^{c}{}_{ac}~{}.$ (49)
@@
$K_{bcd}=-K_{bdc}~{}.$ (50)
The projective Cotton-York tensor $K_{bcd}$ is clearly a tensor on both $\mathcal{M}$ and $\mathcal{N}$ as well as projectively invariant. In Eq. (46), the first line demonstrates the tensor character on $\mathcal{N}$ as recall both $g_{\alpha}$ and $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}$ are tensors on $\mathcal{N}$. The second line demonstrates projective invariance even when restricted to the manifold $\mathcal{M}$ as each of $\mathcal{D}_{ab}$, $g_{a}$, and $\mathcal{K}^{a}{}_{bcd}$ are projectively invariant. The third line demonstrates the tensor character on the manifold $\mathcal{M}$ as each of $\mathcal{P}_{ab}$, ¢$\text{\textcent}_{a}$, and $\mathcal{K}^{a}{}_{bcd}$ are tensors on $\mathcal{M}$. None of the lines alone demonstrate both tensorality and projective invariance on the manifold $\mathcal{M}$, once again showing the importance of defining the covariant and projectively invariant ${\bar{{\mathcal{P}}}}_{ab}$ and ¯$\bar{\nabla}_{a}$ that we will introduce later.
The TW action is the sum of a projective Einstein-Hilbert action and a projective Gauss-Bonnet action
@@
[…]
Our goal in this paper is to maintain this covariance and manifest projective invariance after reducing to the ${\rm d}$-dimensional manifold $\mathcal{M}$ [^2] Here and throughout, by manifest projective invariance we mean that each object, such as ${\mathcal{K}}_{abcd}$ in the action, is projectively invariant.. To illuminate the behavior on the spacetime manifold $\mathcal{M}$ we can expand $S$ and integrate out $\lambda$
$S=$ $\int d^{\rm d}x\sqrt{|g|}\mathcal{L}$ (53)
$\mathcal{L}=$ $-\frac{1}{2\kappa_{0}}(\mathcal{K}+2\Lambda_{0})+\mathbb{L}$ (54)
$\mathbb{L}=$ $J_{0}c\left[\lambda_{0}^{2}K_{bcd}K^{bcd}-\mathcal{K}_{abcd}\mathcal{K}^{abcd}+4\mathcal{K}_{bd}\mathcal{K}^{bd}-\mathcal{K}^{2}\right].$ (55)
The full TW Lagrangian is denoted by $\mathcal{L}$ that includes the projective Gauss-Bonnet (PGB) Lagrangian denoted by $\mathbb{L}$. The rescaled coupling constants are
@@
On $\mathcal{M}$ we use the notation that indices are raised and lowered with $g^{ab}$ and $g_{ab}$, respectively
${\mathcal{K}}_{abcd}\equiv$ $g_{ma}{\mathcal{K}}^{m}{}_{bcd}$ (57)
${\mathcal{K}}^{abcd}=$ $g^{bp}g^{cm}g^{dn}{\mathcal{K}}^{a}{}_{pmn}$ (58)
$K^{bcd}\equiv$ $g^{bp}g^{cm}g^{dn}K_{pmn}~{}.$ (59)
@@
On $\mathcal{N}$ whether indices are raised with $g^{ab}$ or $G^{\alpha\beta}$ is unambiguous for the projective Cotton-York tensor and the last three indices of the projective Riemann tensor $\mathcal{K}^{a}{}_{bcd}$. This is because $G^{ab}=g^{ab}$ and the lambda components of these tensor vanish aside from $\mathcal{K}^{\lambda}{}_{bcd}$. Care must be taken, however, to be clear what is meant by lowering the first index of $\mathcal{K}^{a}{}_{bcd}$. On $\mathcal{M}$ we no longer have the full $G_{\alpha\beta}$ as its $\lambda$ components have been integrated out, so raising and lowering is done via $g^{ab}$ and $g_{ab}$, respectively.
@@
Quadratic dynamics for the fields can be seen in the projective Cotton-York tensor term when expanded using Eq. (46). The curvature tensors contain interactions as can be seen upon expanding in terms of the ${\rm d}$-dimensional $\mathcal{D}_{ab}$ and $\Pi^{c}{}_{ab}$ or ${\mathcal{P}}_{ab}$ and $\Gamma^{c}{}_{ab}$ as
$\mathcal{K}^{a}_{\ bcd}=$ $\mathcal{R}^{a}_{\ bcd}+\delta^{a}_{\ [c}\mathcal{D}_{d]b}$ (60)
$=$ $R^{a}_{\ bcd}+\delta^{a}_{\ [c}\mathcal{P}_{d]b}-\delta^{a}_{\ b}\mathcal{P}_{[cd]}$ (61)
$\mathcal{K}_{mn}=$ $\mathcal{R}_{mn}+({\rm d}-1)\mathcal{D}_{mn}$ (62)
$=$ $R_{mn}+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}\mathcal{P}_{[mn]}$ (63)
$\mathcal{K}=$ $\mathcal{R}+({\rm d}-1)\mathcal{D}=R+({\rm d}-1)\mathcal{P}~{}~{}~{}$ (64)
$\mathcal{R}\equiv$ $g^{ab}\mathcal{R}_{ab}~{},~{}\mathcal{D}\equiv g^{ab}\mathcal{D}_{ab}$ (65)
$R\equiv$ $g^{ab}R_{ab}~{},~{}{\mathcal{P}}\equiv g^{ab}{\mathcal{P}}_{ab}~{}.$ (66)
@@
All other curvature components vanish aside from $\mathcal{K}^{\lambda}{}_{bcd}$ which can be solved for using the expressions of the projective Cotton-York tensor in Eq. (46). We see explicitly each object in TW gravity is not covariant and manifestly projective invariant on $\mathcal{M}$ when written in terms of these expansions. If the curvature tensors are expanded through the first equality in each of the above, they are manifestly projective invariant, but not covariant. If the second equality is used in each, they are no longer manifestly projective invariant, but they are covariant. The same is true for the projective Cotton-York tensor as was shown in Eq. (46). In the next section, we introduce the covariant and manifestly projective invariant ${\bar{{\mathcal{P}}}}_{ab}$ and ¯$\bar{\nabla}_{a}$ and expand the above tensors in terms of these.
@@
We conclude this section with a brief aside on the values of the physical constants of TW gravity that are consistent with physical phenomena, specifically early Universe inflation. For ${\rm d}=4$ the units are
$\begin{split}[J_{0}]=\frac{ML^{2}}{T}~{}~{}~{},~{}~{}~{}[\mathcal{D}_{ab}]&=[R_{ab}]=L^{-2}\cr\left[\kappa_{0}\right]=\frac{T^{2}}{ML}~{}.\end{split}$ (67)
@@
$M_{p}=\sqrt{\frac{\hbar c}{8\pi G}}$ (68)
@@
a recent analysis in [24] suggests that if $\lambda_{0}\lesssim 6\cdot 10^{6}M_{p}^{-1}$ the inflaton $\phi\propto{\mathcal{P}}$ appearing naturally in TW gravity would have provided early Universe inflation matching current inflationary cosmology data within the $95\%$ confidence level. Specifically, it was found that to produce $N=60$ e-folds, the parameters would have to satisfy the constraints $\lambda_{0}^{2}\lesssim 10^{3}J_{0}\kappa_{0}$ and $10^{8}\lesssim J_{0}\lesssim 4\cdot 10^{10}$ all within the $95\%$ confidence level. Furthermore, a value of $\kappa_{0}\sim M_{p}^{-2}$ would result in Einstein-Hilbert gravity + cosmological constant at the end of inflation within the $95\%$ confidence level  [^3] The value of $\kappa_{0}$ was not discussed in this way in [24]. However, it is easy to see that the TW action reduces to $(-2\kappa_{0})^{-1}(1+8J_{0}\kappa_{0}\lambda_{0}^{-1}\phi)R$ plus cosmological constant at the end of inflation if $\phi\propto{\mathcal{P}}$ settles to a constant value, with the traceless parts of $\mathcal{P}_{ab}$ equal to zero and assuming a Levi-Civita connection. The analysis in [24] leads to $18J_{0}\kappa_{0}\lambda_{0}^{-1}\phi<<$ 1 at the end of $N=60$ e-folds of inflation within the $95\%$ confidence level. Thus the TW action reduces to approximately $(-2\kappa_{0})^{-1}R$ plus cosmological constant at the end of inflation, where we identify $\kappa_{0}\sim M_{p}^{-2}$ to match general relativity.. Putting this all together results in the above limit on $\lambda_{0}$. This suggest that the characteristic projective scale defined by $\lambda_{0}$ is less than a few million Planck lengths if it is TW gravity itself that is responsible for inflation. The importance of these results is that it appears inflation is automatically is built into the geometry of TW gravity with no additional field needed. The above analysis for the limit on $\lambda_{0}$ relies on the assumption that TW gravity must precisely match Einstein-Hilbert gravity + cosmological constant at the end of inflation to then transition to radiation dominance. Whether or not that is true depends on the specific interactions of TW gravity with matter because matter acts like radiation in the early Universe. These interactions are an ongoing avenue of research [18, 19, 21, 20].
II.4 Manifestly Projective Invariant form of the TW Field Equations